If someone invested 13000.00 and gets 12% interest, how much more money does she need in an acct that pays 11% interest so the average rate of return on the investments in 11.667%?
Answer by mananth(16946)

percent Principal $
12.00 13000
11.00 x
11.67 13000 + x

12 * 13000 + 11 x = 11.67 * ( 13000 + x )
156000 + 11 x = 11.67 * 13000 + 11.67 x
156000 + 11 x = 151671 + 11.67 x
11 x - 11.67 x = 151671 - 156000
-0.67 x = -4329
/ -0.67
x = 6490.25
m.ananth@hotmail.ca